# From Query to Logic: Ontology-Driven Multi-Hop Reasoning in LLMs

### v1

@@ Introduction
Large Language Models have demonstrated significant success in knowledge-based question answering (Brown et al. 2020; DeepSeek-AI 2025; Yang et al. 2025). However, they continue to face challenges in MQA tasks (Wang et al. 2024; Shao et al. 2023a), which require integrating and reasoning over multiple, discrete sources of information. A significant challenge, as highlighted in Ju et al. (2024), is that LLMs tend to rely on guessing derived from training data rather than performing reasoning in MQA tasks. Therefore, the key lies in advancing beyond the generation of factually correct chains toward empowering LLMs to uncover and leverage deeper conceptual relationships between entities—a capability essential for true multi-hop understanding.
@@
For instance, as shown in Figure 1, given the question $Q$ = “What government followed the monarch who retranslated the Reflections into French from the country that allied with America after the Battle of Saratoga?”, this stage extracts:
  • Classes and Entities ($C_{Q}$): A set of type assignments such as Country(America,c), Book(TheReflections), Monarch(m), Government(ans) and Battle(Saratoga), where c, m, and ans are a unknown variables.
  • Relations ($R_{Q}$): A set of predicates including retranslated, rulerOf, followed, finished and alliedWith.
@@
As illustrated in Figure 1, the question $Q$ is transformed into the following formula $\Phi$, accompanied by type declarations. The predicate names in the example are chosen for clarity.
Type Constraints: type(m,Monarch), type(c,Country), type(ans,Government)
@@
[…]
@@
Input: Input question $Q$, Large Language Model $\mathcal{LLM}$, Retriever Engine $\mathcal{E}_{Retriever}$, Ontology Prompt $\mathcal{P}_{Onto}$, FOL Prompt $\mathcal{P}_{FOL}$, Decomposition Prompt $\mathcal{P}_{Decomp}$
@@
9:for $i=1$ to $n$ do
10:  $Q^{\prime}_{i}\leftarrow{substitute}(Q_{i},A_{ans})$
11$\triangleright$ Substitute placeholders
12:  $ans_{i}\leftarrow\mathcal{LLM}([\mathcal{E}_{Retriever}(Q^{\prime}_{i})\oplus Q^{\prime}_{i}])$
13$\triangleright$ Execute the sub-question
14:  Append $ans_{i}$ to $A_{ans}$
15:end for
@@
Unless specified otherwise, all experiments utilized the gpt-3.5-turbo API [^2] Developed by OpenAI, a specific model version is gpt-3.5-turbo-0125. as the base LLM. For non-retrieval baselines (e.g., NoCoT, CoT, and RAG), models were prompted to enclose final answers in <answer> tags to standardize evaluation. The NoCoT (DeepSeek-R1) baseline adopted the same prompt structure but used the DeepSeek-R1 model [^3] Developed by DeepSeek, a specific model version is DeepSeek-R1-0528.. For all retrieval-based methods (ReAct, LPKG, and our proposed ORACLE), we implemented a unified retriever module. The ReAct baseline integrated this retriever with gpt-3.5-turbo as its core agent. Similarly, LPKG employed gpt-3.5-turbo to perform its in-context learning-based planning. As a method designed for powerful, general-purpose models, LPKG provides a strong point of comparison, which contrasts with our training-free framework. Detailed implementation can be found in the Appendix.
@@
Our analysis reveals two critical findings. First, ORACLE consistently produces higher-quality reasoning paths. As shown in Figure 4, ORACLE achieves a higher average Reasoning $F_{1}$ score than LPKG across all question types, with an overall score of 0.648 compared to LPKG’s 0.471. This indicates that ORACLE’s planning module is fundamentally more effective. Figure 3 reinforces this by showing the respective proportion of high- and low-quality paths. For every question type, ORACLE generates a larger fraction of high-quality paths. For example, on Compositional questions, 90.83% of ORACLE’s reasoning paths exceed the $F_{1}>0.5$ threshold, whereas only 63.76% of LPKG’s paths do.
Figure 5: Impact of Intermediate Reasoning Quality. The analysis is presented for both the LPKG and ORACLE methods across the four reasoning categories. The line color signifies the correlation between reasoning quality and final answer quality for that category: green indicates a positive correlation, while red indicates a weaker or negative correlation.
@@
As illustrated in Figure 6, our method’s decomposition aligns closely with the ground truth. A perfect match (+0+0+ 0 deviation) is the most frequent outcome, occurring in 46.40% of cases. Overall, 98.00% of generated plans deviate by at most one step (±1$\pm$ 1± 1), with 39.20% over-decomposing by one (+1+1+ 1) and 12.40% under-decomposing by one (1-1- 1).
Notably, the model’s performance, measured by the F1 score, peaks at 0.686 for a +1+1+ 1 deviation. This score is substantially higher than the F1 score of 0.479 achieved for a perfect match (+0+0+ 0 deviation). This suggests a clear asymmetry in reasoning errors: including a redundant step is far less detrimental, and often beneficial, than omitting a crucial one. Conversely, under-decomposition (1-1- 1 deviation) causes a drop in performance to an F1 score of 0.408, indicating that failing to generate a required sub-question impairs the reasoning process.
@@ Appendix A Appendix
(2) Lookup[keyword], which returns the next sentence containing keyword in the current passage.
@@ 4. Logical Interpretation: Explain the meaning of the logical expression in natural language
Given the scarcity of research evaluating the internal reasoning processes of LLMs, and substantial evidence showing that models can produce correct answers via fallacious reasoning, a more granular assessment is necessary. As noted by Ju et al. (2024), when faced with a query such as, “On which continent is the home country of the 2022 FIFA World Cup winner?,” an LLM might not follow the correct reasoning chain, “Winner $\rightarrow$ Argentina $\rightarrow$ South America.” Instead, it may leverage a shortcut learned during training, directly linking “2022 World Cup Winner” to “South America.”
@@
Table 5: A comparison of NoCoT and CoT
@@
For our evaluation, we transform the decomposed sub-questions and their answers from both ORACLE and LPKG into the same concatenated string of triples. We then measure the fidelity of each method’s reasoning by computing the F1 score between the generated string and the ground-truth string. In a parallel analysis, we also quantify the deviation between the number of sub-questions our methods generate versus the number provided in the ground-truth decomposition, allowing us to measure how the granularity of question decomposition affects task performance. For a detailed implementation, please refer to the accompanying code.
Cases Study

### v2

@@ Introduction
Large Language Models have demonstrated significant success in knowledge-based question answering (Brown et al. 2020; DeepSeek-AI 2025; Yang et al. 2025). However, they continue to face challenges in MQA tasks (Wang et al. 2024; Shao et al. 2023a), which require integrating and reasoning over multiple, discrete sources of information. A significant challenge, as highlighted in Ju et al. (2024), is that LLMs tend to rely on guessing derived from training data rather than actually reasoning in MQA tasks. Therefore, the key lies in advancing beyond the generation of factually correct chains toward empowering LLMs to uncover and leverage deeper conceptual relationships between entities—a capability essential for true multi-hop understanding.
@@
  • Prompt Concatenation ($\oplus$): An operator used to denote the sequential concatenation of prompts or text segments, i.e., $P_{1}\oplus P_{2}$ appends $P_{2}$ after $P_{1}$ with a delimiter (such as a newline or separator token).
@@
For instance, as illustrated in Figure 1, given the question $Q$ = “What government followed the monarch who retranslated the Reflections into French from the country that allied with America after the Battle of Saratoga?”, this stage extracts:
  • Classes and Entities ($C_{Q}$):
    $\text{type}(\text{The Reflections},Book),$
    $\text{type}(m,Monarch),$
    $\text{type}(c,Country),$
    $\text{type}(Saratoga,Battle),$
    $\text{type}(ans,Government)$
    where $m$, $c$, and ansans are variables. Notably, ansans denotes the final answer entity (the Government).
  • Relations ($R_{Q}$): $m\xrightarrow{\texttt{retranslated}}$ The Reflections; $m\xrightarrow{\texttt{rulerOf}}c$; $c\xrightarrow{\texttt{followed}}ans$; $c\xrightarrow{\texttt{alliedWith}}$ America; Saratoga $\xrightarrow{\texttt{finished}}$ America.
This structure corresponds to the ontology graph shown in Figure 1, where nodes denote classes and instantiated entities, and directed edges denote relations among them.
Why Ontology is Necessary.
While LLMs inherently possess background knowledge about many entities, direct decoding often relies on local token-level similarity, which can easily introduce irrelevant entities into the reasoning chain. For instance, as illustrated in our case study (Appendix A), a baseline decomposition mistakenly aligns The Blonde from Singapore with The Blonde from Peking, a phenomenon we term logicalized semantic drift.
To mitigate this issue, we introduce a lightweight ontology as an intermediate scaffold. The ontology enforces (i) type constraints, ensuring that variables are grounded to the correct conceptual classes, (ii) relational consistency, aligning predicates with the intended semantic hierarchy, and (iii) path stability, as demonstrated in the appendix case studies (Appendix A). Together, these properties substantially reduce semantic drift and provide a global structural perspective before entering the FOL construction stage.
@@
As illustrated in Figure 1, the question $Q$ is transformed into the following formula $\Phi$, accompanied by type declarations. The predicate names in the example are aligned with the ontology graph for clarity.
Type Constraints:
$\text{type}(m,Monarch),$
$\text{type}(c,Country),$
$\text{type}(ans,Government)$
@@
$\Phi=$ $\exists~ans,m,c\quad$
$\land\text{retranslated}(m,\text{The Reflections},French)$
$\land\text{rulerOf}(m,c)$
$\land\text{alliedWith}(c,America)$
$\land\text{finished}(Saratoga,America)$
$\land\text{followed}(c,ans)$
@@
Input: Input question $Q$, Large Language Model $\mathcal{LLM}$, Retriever $\mathcal{R}$, Ontology Prompt $\mathcal{P}_{Onto}$, FOL Prompt $\mathcal{P}_{FOL}$, Decomposition Prompt $\mathcal{P}_{Decomp}$
@@
5:for $i=1$ to $n$ do
6:  $Q^{\prime}_{i}\leftarrow{substitute}(Q_{i},A_{ans})$
7:  $ans_{i}\leftarrow\mathcal{LLM}([\mathcal{R}(Q^{\prime}_{i})\oplus Q^{\prime}_{i}])$
8:  Append $ans_{i}$ to $A_{ans}$
9:end for
@@
Unless specified otherwise, all experiments utilized the gpt-3.5-turbo API [^2] Developed by OpenAI; the specific model version we use is gpt-3.5-turbo-0125. as the base LLM. For non-retrieval baselines (e.g., NoCoT, CoT, and RAG), models were prompted to enclose final answers in <answer> tags to standardize evaluation. The NoCoT (DeepSeek-R1) baseline adopted the same prompt structure but used the DeepSeek-R1 model [^3] Developed by DeepSeek; the specific model version we use is DeepSeek-R1-0528.. For all retrieval-based methods (ReAct, LPKG, and our proposed ORACLE), we implemented a unified retriever module. The ReAct baseline integrated this retriever with gpt-3.5-turbo as its core agent. Similarly, LPKG employed gpt-3.5-turbo to perform its in-context learning-based planning. As a method designed for powerful, general-purpose models, LPKG provides a strong point of comparison, which contrasts with our training-free framework. Detailed implementation can be found in the Appendix.
@@
Our analysis reveals two critical findings. First, ORACLE consistently produces higher-quality reasoning paths. As shown in Figure 4, ORACLE achieves a higher average Reasoning $F_{1}$ score than LPKG across all question types, with an overall score of 0.648 compared to LPKG’s 0.611. This indicates that ORACLE’s planning module is fundamentally more effective. Figure 3 reinforces this by showing the respective proportion of high- and low-quality paths. For every question type, ORACLE generates a larger fraction of high-quality paths. For example, on Compositional questions, 90.83% of ORACLE’s reasoning paths exceed the $F_{1}>0.5$ threshold, whereas only 63.76% of LPKG’s paths do.
Figure 5: Impact of Intermediate Reasoning Quality. The Analysis Is Presented for Both the LPKG and ORACLE Methods Across the Four Reasoning Categories. The Line Color Indicates the Correlation Between Reasoning Quality and Final Answer Quality for That Category: Green Indicates a Positive Correlation, While Red Indicates a Weaker or Negative Correlation.
@@
As illustrated in Figure 6, our method’s decomposition aligns closely with the ground truth. A perfect match (+0+0 deviation) is the most frequent outcome, occurring in 46.40% of cases. Overall, 98.00% of generated plans deviate by at most one step ($\pm 1$), with 39.20% over-decomposing by one (+1+1) and 12.40% under-decomposing by one (1-1).
Notably, the model’s performance, measured by the F1 score, peaks at 0.686 for a +1+1 deviation. This score is substantially higher than the F1 score of 0.479 achieved for a perfect match (+0+0 deviation). This suggests a clear asymmetry in reasoning errors: including a redundant step is far less detrimental, and often beneficial, than omitting a crucial one. Conversely, under-decomposition (1-1 deviation) causes a drop in performance to an F1 score of 0.408, indicating that failing to generate a required sub-question impairs the reasoning process.
More experiments across Model Sizes on Qwen2.5
Firstly, we note that we did not extend validation to dedicated reasoning models, since without fine-tuning, such models typically struggle with complex instruction-following. Thus, our experiments focus on general-purpose models where ontology-guided reasoning can be more faithfully evaluated.
To further validate the effectiveness of our method beyond the main setting, we evaluate LPKG and ORACLE on a series of models, including Qwen2.5 with 7B, 14B, 32B, and 72B parameters. The results are summarized in Table 3, covering three metrics: Exact Match (EM), F1, and Reasoning F1.
Model Method EM F1 Reason F1
GPT-3.5 LPKG 0.379 0.452 0.611
ORACLE 0.469 0.548 0.650
Qwen2.5-7B LPKG 0.432 0.533 0.620
ORACLE 0.447 0.507 0.598
Qwen2.5-14B LPKG 0.471 0.571 0.637
ORACLE 0.494 0.587 0.654
Qwen2.5-32B LPKG 0.440 0.531 0.665
ORACLE 0.527 0.635 0.679
Qwen2.5-72B LPKG 0.485 0.584 0.659
ORACLE 0.557 0.638 0.688
Table 3: Vertical Comparison of LPKG and ORACLE Across Different Models. Metrics Include Exact Match (EM), F1, and Reasoning F1. The Superscript $\star$ Indicates the Only Case Where LPKG Outperforms ORACLE in Reasoning F1.
From the results, we observe the following:
  • Overall Superiority of ORACLE. Except for the 7B model, ORACLE consistently outperforms LPKG across all metrics and model sizes. This validates that our ontology-driven approach yields stable improvements with larger and more capable models.
  • 7B Exception. On Qwen2.5-7B, LPKG achieves higher Reasoning F1 (0.6202 vs. 0.5983). We hypothesize that the limited parameter capacity restricts the model’s ability to leverage ontology guidance effectively. In contrast, LPKG benefits more from its detailed prompt design, which encodes more external knowledge rather than relying on the model’s internal reasoning.
  • Scaling Effects. As parameter size increases (32B vs. 72B), F1 scores become similar, but Reasoning F1 continues to improve (0.6792 $\rightarrow$ 0.6879). This suggests that larger models not only strengthen internal knowledge representation but also enhance their ability to integrate information for reasoning.
  • Architecture Efficiency. Comparing GPT-3.5 (175B) with Qwen2.5-14B, we find comparable or even superior performance from the smaller Qwen2.5 model, indicating that newer architectures deliver higher parameter efficiency. This trend supports the view that sparsity and architectural advances are as critical as sheer scale.
@@ Appendix A Appendix
(2) Lookup[keyword], which returns the next sentence containing the keyword in the current passage.
@@ 4. Logical Interpretation: Explain the meaning of the logical expression in natural language
Given the scarcity of research evaluating the internal reasoning processes of LLMs, and substantial evidence showing that models can produce correct answers via fallacious reasoning, a more granular assessment is necessary. As noted by Ju et al. (2024), when faced with a query such as, “On which continent is the home country of the 2022 FIFA World Cup winner?” an LLM might not follow the correct reasoning chain, “Winner $\rightarrow$ Argentina $\rightarrow$ South America.” Instead, it may leverage a shortcut learned during training, directly linking “2022 World Cup Winner” to “South America.”
@@
Table 6: Comparison of NoCoT and CoT
@@
For our evaluation, we transform the decomposed sub-questions and their answers from both ORACLE and LPKG into the same concatenated string of triples. To ensure fair comparison with the gold reasoning path, we first apply a preprocessing step to the generated reasoning text, including lowercasing, stopword removal, and lemmatization. This yields a normalized representation of the reasoning process that contains only the essential entities and relations. We then align this representation with the ground-truth reasoning path provided in the dataset, also expressed as a concatenated sequence of triples.
Formally, let $G=\{g_{1},g_{2},\dots,g_{m}\}$ denote the set of tokens (entities and relations) in the ground-truth reasoning path, and $P_{\text{gen}}=\{p_{1},p_{2},\dots,p_{n}\}$ denote the tokens extracted from the model-generated reasoning path after preprocessing. We compute precision $P$ and recall $R$ as:
$P=\frac{|G\cap P_{\text{gen}}|}{|P_{\text{gen}}|},\quad R=\frac{|G\cap P_{\text{gen}}|}{|G|}$
and define the Reasoning F1 score as:
$\text{ReasonF1}=\frac{2\cdot P\cdot R}{P+R}.$
In parallel, we also compute the sub-question count deviation to measure the granularity of decomposition. Let $k^{*}$ denote the ground-truth number of reasoning steps and $k_{\text{gen}}$ the number generated by the model. The deviation is defined as:
$\Delta k=k_{\text{gen}}-k^{*}.$
This enables us to analyze how over-decomposition or under-decomposition impacts task performance.
Case Studies
@@
LPKG and ORACLE
To further illustrate the advantages of our ontology-driven decomposition, we present two representative case studies comparing ORACLE with the LPKG baseline.
LPKG ORACLE
Question Which film has the director who died first, The Piper’s Price or The Blonde From Singapore?
Reasoning Process Which film is directed by the director of The Piper? $\rightarrow$ The Faded Woman Which film is directed by the director of The Blonde From Singapore? $\rightarrow$ The Blonde from Peking Who directed The Piper’s Price? $\rightarrow$ Joe De Grasse Who directed The Blonde From Singapore? $\rightarrow$ Edward Dmytryk When did [Joe De Grasse] die? $\rightarrow$ May 25, 1940 When did [Edward Dmytryk] die? $\rightarrow$ July 1, 1999 Compare death dates $\rightarrow$ May 25, 1940
Ground Truth Path The Piper’s Price $\rightarrow$ Joe De Grasse $\rightarrow$ May 25, 1940; The Blonde From Singapore $\rightarrow$ Edward Dmytryk $\rightarrow$ July 1, 1999
ReasonF1 0.30 0.66
Table 7: Case study comparing LPKG and ORACLE on a film director comparison question.
The first case (Table 7) is a comparison-type query asking which film’s director died first. LPKG mistakenly interprets the sub-question as “find another film directed by the same director,” which leads to irrelevant reasoning paths. In contrast, ORACLE first extracts the ontology to anchor the key entities (films, directors, and death dates), then generates a reasoning chain that matches the ground truth. This results in a substantially higher ReasonF1, indicating closer alignment with the gold reasoning path.
LPKG ORACLE
Question Where was the father of Marianus V of Arborea born?
Reasoning Process Who is Marianus V of Arborea? $\rightarrow$ Marianus V was the Judge of Arborea Where was the father of “Marianus V was the Judge of Arborea” born? $\rightarrow$ Castel Genovese Who is the father of Marianus V of Arborea? $\rightarrow$ Brancaleone Doria Where was [Brancaleone Doria] born? $\rightarrow$ Republic of Genoa
Ground Truth Path Marianus V of Arborea $\rightarrow$ father $\rightarrow$ Brancaleone Doria $\rightarrow$ place of birth $\rightarrow$ Sardinia
ReasonF1 0.27 0.70
Table 8: Case study comparing LPKG and ORACLE on a familial relation query.
The second case (Table 8) is a relation-based query about the birthplace of Marianus V of Arborea’s father. Here, LPKG drifts to a descriptive path focusing on Marianus V himself, while ORACLE correctly identifies the parent entity (Brancaleone Doria) and queries his birthplace. Although minor deviations from the ground answer exist, the ontology-driven reasoning still yields a reasoning chain much closer to the ground truth, again reflected in a higher ReasonF1 score.
Together, these examples demonstrate that ORACLE not only avoids semantic drift in sub-question decomposition but also consistently produces reasoning paths that are more faithful to the intended logical structure, leading to more interpretable and reliable answers across diverse query types. Furthermore, we observe that across multiple runs, ORACLE tends to generate reasoning processes that are relatively stable, a property largely attributable to the guidance of ontology extraction at the planning stage.